OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#4296 — DEC-iX
Scheduled Maintenance Report for Network & Infrastructure
Completed
All links towards the DEC-IX are down. More information are forthcoming.

Update(s):

Date: 2010-06-18 06:20:06 UTC
dec-ix faced some difficulties but bgp sessions are now forwarded.

Update 2010-06-18 05:35 UTC

We still have problems with sw021, all LACP port channels and some single ports did not
come up after the reboot and the switch forgot parts of the configuration. We restored
configuration and rebooted again, if this does not work we will revert to the old software.

Date: 2010-06-18 06:18:48 UTC
The physical links are now reset but we are still isolated from lan peering at the moment.

Date: 2010-06-18 06:16:31 UTC
The corruption is due to an urgent maintenance of the switch on which we are linked to DEC-IX. The service has to be reset in ten minutes.
Posted Jun 18, 2010 - 06:06 UTC